An ice cream company collected data on their ice cream cones sales over a month in July in a Chicago suburb, along with daily temperature and the weather. The company is interested to develop a correlation between ice cream sales to the hot weather. Market research showed that more people come out in certain neighborhoods, to either enjoy the nice weather, or venture out if they do not have air conditioning in their apartments. The Chicago Police also tracked crime statistics during the same period. Crime statistics included murder, assault, robbery, battery, burglary, theft and motor vehicle theft. The data are shown below:

July

Day Temp (F)

Weather

Ice cream sales (units)

Crime stats reported

1

83

Thunderstorm

590

201

2

81

Thunderstorm

610

220

3

84

Thunderstorm

640

199

4

79

Partly sunny

490

195

5

80

Mostly sunny

550

187

6

84

Sunshine

710

280

7

84

Sunshine

690

261

8

86

Thunderstorm

750

310

9

83

Shower

720

254

10

86

Partly sunny

850

300

11

83

Partly sunny

690

219

12

84

Cloudy

750

275

13

81

Thunderstorm

450

156

14

82

Thunderstorm

550

210

15

80

Heavy rain

25

98

16

81

Heavy rain

78

110

17

86

Sunshine

790

256

18

81

Sunshine

530

145

19

81

Sunshine

490

199

20

80

Sunshine

620

245

21

80

Sunshine

690

260

22

79

Sunshine

540

159

23

81

Partly sunny

610

299

24

80

Partly sunny

590

239

25

81

Partly sunny

590

250

26

80

Sunshine

580

200

27

87

Sunshine

880

300

28

91

Sunshine

1,059

361

29

90

Sunshine

1,000

401

30

91

Partly sunny

960

375

31

88

Partly sunny

890

360

                           [ Select ]                       ["-3892.2 + 40.1(x); r^2 = .78", "-3462.4 + 49.4(x); r^2 = .78", "-2362.5 + 39.2(x); r^2 = 0", "-3432.6 + 41.3(x); r^2 = .61"]         Develop a linear regression model for ice cream sales over daily temperature. Show the linear equation in the form of y = ax + b, and the correlation of determination (r^2).

                           [ Select ]                       ["1201", "1101", "1001", "1181"]         What would be the projected forecast of ice cream sales in units, for daily temperature of 94 F?

                           [ Select ]                       ["-2808.1 + 41.9(x); r^2 = .86", "-1362.5 + 33.2(x); r^2 = .67", "-3932.6 + 51.3(x); r^2 = .93", "-2892.2 + 60.1(x); r^2 = .55"]         On July 15 & 16 there were heavy down pour of rain, which might have prevented some to venture out to purchase ice cream during the day. If you were to override those 2 data points, what would be the linear regression model be (by deleting July 15 & 16 data).

                           [ Select ]                       ["second correlation is a better forecast", "need more data", "no different", "first correlation is a better forecast"]         Compare the two correlation coefficients, which would be considered a better forecast for ice cream sales

                           [ Select ]                       ["92.2 + 10.1(x); r^2 = .61", "50.1 + 11.8(x); r^2 = .96", "32.6 + .39(x); r^2 = .71", "46.8 + .30(x); r^2 = .82"]         Develop a linear regression on ice cream sales to crime statistics. Show the linear equation in the form of y = ax + b, and the correlation of determination (r^2).

                           [ Select ]                       ["no, correlation does not imply causality", "yes, strong correlation does imply causality"]         Does this correlation demonstrate causation, that high ice cream sales cause crime statistics to go up?
Answer #1

The output is shown below

The linear regression model is Y = -3462.45 + 49.39X and the correlation coefficient is 0.77. Thus the correct answer will be "-3462.4 + 49.4(x); r^2 = .78"

Projected forecast with when X = 94 will be -3462.4 + 49.4*94 = 1181.2. Thus the correct answer will be 1181.

After deleting July 15 and 16 the output is shown below.

The linear regression model is Y = -2808.06 + 41.92X and the coefficient of determination is 0.85. Thus the correct answer will be "-2808.1 + 41.9(x); r^2 = .86"

Comparing the correlation coefficients, the correct answer will be

Second correlation is a better forecast

The output is show below

The regression model is Y = 46.78 + 0.3X and coefficient of determination is 0.81. Thus the correct answer will be "46.8 + .30(x); r^2 = .82"

No. Correlation does not imply causality